OpenStudy (anonymous):

Separate 65 into two parts so that 3 times the larger will be 6 more than 6 times the smaller. Separate 65 into two parts so that 3 times the larger will be 6 more than 6 times the smaller. @Mathematics

OpenStudy (anonymous):

44

OpenStudy (anonymous):

could you tell me how its 44 do i could understand please

OpenStudy (anonymous):

let the larger number be x 3x - 6 = 6*(65-x) now solve for x!!

OpenStudy (anonymous):

396/9= 43.33 or 43